# MELTing point: Mobile Evaluation of Language Transformers

### v2

@@ Abstract.
Transformers have revolutionized the machine learning landscape, gradually making their way into everyday tasks and equipping our computers with “sparks of intelligence”. However, their runtime requirements have prevented them from being broadly deployed on mobile. As personal devices become increasingly powerful at the consumer edge and prompt privacy becomes an ever more pressing issue, we explore the current state of mobile execution of Large Language Models (LLMs). To achieve this, we have created our own automation infrastructure, MELT, which supports the headless execution and benchmarking of LLMs on device, supporting different models, devices and frameworks, including Android, iOS and Nvidia Jetson devices. We evaluate popular instruction fine-tuned LLMs and leverage different frameworks to measure their end-to-end and granular performance, tracing their memory and energy requirements along the way.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Lately, however, transformers (Vaswani et al., 2017) have become the go-to architecture for deep learning models, with attention mechanisms offering unparalleled performance and the ability to model long sequence data with fewer inductive biases. Applied across modalities, including vision, speech and text (Dosovitskiy et al., 2021; Radford et al., 2019, 2023), these models have demonstrated significant performance benefits, across modeling and generation tasks. Another key benefit of these models has been their ability to scale to very large sizes, both in terms of data ingestion and parameter size, without their performance plateauing (Touvron et al., 2023b). This has given birth to “foundation models”, large models that are trained on large corpora of data and act as universal backbones for a series of downstream tasks. For example, Large Language Models (LLMs) (Touvron et al., 2023b; Chung et al., 2022; Hoffmann et al., 2022) have been trained on large parts of the Internet (commoncrawl.org, 2024; Raffel et al., 2020) and are able to tackle downstream tasks without explicit training (Wang et al., 2023; Dong et al., 2023; Vernikos et al., 2023).
@@
To this end, we have created our own infrastructure, named MELT  [^1] We plan to release soon the design and software of our infrastructure to facilitate reproducibility and democratize on-device deployment of LLMs., designed to interact, trace and benchmark LLMs across ML frameworks, devices, and ecosystems. With our tool, we automate the interaction with instruction fine-tuned models and capture events and metrics of interest at a granular level, both in terms of performance as well as energy. While in this paper MELT is deployed for LLMs inference, our infrastructure is general enough to support the tracing of any workload, without user intervention. To the best of our knowledge, our tool is the first to support granular on-device energy measurements across device targets (i.e., Android, iOS, Linux) with realistic interactions.
Research Questions. Given the constant developments in mobile and embedded System-On-Chips (SoCs) and the meteoric rise of LLMs, we aspire to measure the deployability of Large Language Models at the consumer edge and identify the bottlenecks that prevent the broad deployment of such workloads. Specifically, the research questions that we aim to answer with this study are the following:
@@ 2.1. Transformer Preliminaries
The main contribution of transformers has undoubtedly been the attention mechanism, which captures the relationship between tokens in a sequence from a single source (self-attention) or multiple sources (multi-head attention). Attention is calculated as $A(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$, where $Q$, $K$, $V$ represent the query, key, and value matrices, respectively and $d_{k}$ the dimensionality of the key matrix. The inner product boosts closer query-key vectors (relevance), softmax normalizes the dot-product and the multiplication with the value results in the relevant value scores being retrieved. The quadratic complexity of attention, with respect to the sequence length (i.e., the prompt or intermediate tokens), is one of the main bottlenecks of deployment, which has given way to alternatives such as sparse (Beltagy et al., 2020) or approximate (Wang et al., 2020; Ainslie et al., 2023) attention mechanisms, as well as attention-free variants as of lately (Gu and Dao, 2023). Context size refers to the maximal window of tokens a transformer block can pay attention to, whereas the maximum generated length refers to the maximum number of tokens generated as output. Generation ends when an <EOS> (end-of-sequence) token is generated. The auto-regressive nature of decoding means that given input sequence $X=\{x_{1},x_{2},\dots,x_{t}\}$, the model generates $x_{t+1}$, which is fed to the next generation step. Key-Value cache (Pope et al., 2023) optimizes this by storing intermediary attention states.
@@ 2.2. Large Language Models
What has made Transformers an instant success has been their applicability to various modalities and their scalability to very large parameter sizes without saturating accuracy (Touvron et al., 2023b). This phenomenon has given birth to Foundation Models (FMs), pretrained on huge corpora of data, i.e., text in our case, and act as a great tool for modeling language and a starting point for fine-tuning on downstream tasks. The task of pretraining usually comprises masked or next-word prediction (self-supervised), whereas downstream tasks can include anything from translation to summarization. Instruction fine-tuning (Ouyang et al., 2022) refers to a specific form of fine-tuning where the model is trained on pairs of input-output instructions. Last, alignment is usually the final step of model tuning, typically through reinforcement learning from human (Ouyang et al., 2022) or automated (Bai et al., 2022) feedback, to promote a certain style or content or reply that “aligns” with values of the creator (e.g., safety). Training cost generally scales down as we move from pretraining downstream, as do data ingestion needs (Zhou et al., 2024).
@@ 2.3. Current State and Motivating Factors
LLMs democratization. Nevertheless, more and more models offer openly their weights, including models from Meta (Touvron et al., 2023a, b), Mistral AI (Jiang et al., 2023), Google (Google Inc., 2024) and Microsoft (Javaheripi, Mojan and Bubeck, Sébastien, 2024). This creates an excellent opportunity for users to deploy their models locally and even personalize them to their preferences, without data ever leaving their device premises. However, such models remain significantly smaller in scale and still require considerable resources to deploy. Towards this end, new frameworks are emerging for enabling local execution of LLMs across different targets (Gerganov, 2023; team, 2023; Hannun et al., 2023; Mark Sherwood, 2024; Alibaba, 2023; tinygrad, 2023). In this effort, quantization (Shen et al., 2020; Frantar et al., 2022; Lin et al., 2023) is one of the most prominent out-of-the-box solutions for reducing their footprint. Yet another enabler towards this democratization is the broad availability of capable SoCs at cost. Indicatively, from our measurements, a recent M2-based Mac Studio can run Llama-2 (Touvron et al., 2023b) 7B model (4-bit quantized) at a sustained 46.8 tokens/sec.
@@ 3.2. JetsonLab
Pseudocode for MELT experiments. Functionality of undefined methods in comment. Prefixed methods run on the device in prefix (e.g., Monsoon, device). Input: PhoneLab, JetsonLab, Monsoon, GPIO, YKUSH, device, $Q^{\text{device}}_{\text{experiments}}$, iterations, samplingFrequency, betweenExpSleep
PowerOn(device) if device.platform == ”ios” :
1       ConnectBT(device) # connect as HID device via Bluetooth UnlockScreen(device) # unlock screen with passcode over HID
2SyncClocks(device) # sync host and guest clocks apiAddress = StartRESTServer() # start REST service on host for exp in $Q^{\text{device}}_{\text{experiments}}$ : # iterate over experiments in the queue
3       Push([exp.model, exp.conversations], device) # push dependencies Apply(exp.conf, exp.model, device) # edit model conf and execution parameters on device for it=0; it¡iterations;++it :
4             StartMonitoring(Monsoon, device) RunExperiment(exp, device) StopMonitoring(Monsoon, device) # disable monitoring CollectMeasurements(exp, device) # get results from FS sleep(betweenExpSleep) # sleep between runs
@@
7             GPIO.EnableRail(device.rails) # enable rail through GPIO YKUSH.PowerOn(device) # enable YKUSH USB of device Monsoon.SetVoutCurr(device) # configure Monsoon power out Wait(device) # wait until device is responsive
8def StartMonitoring(Monsoon, YKUSH, device):
@@
10             YKUSH.DisableUSB() Monsoon.MeasurementMode(”on”, samplingFrequency)
@@
12             Jetsonlab.ScheduleEvents(samplingFrequency) Jetsonlab.Monitor(”on”) # poll SysFS
13def RunExperiment(exp, device, apiAddress):
14       # open app w/ ADB, Bluetooth HID or SSH app = device.OpenApp(exp.backend) Automate(app, model, device) # automate interaction with app http.post(”start”, apiAddress) # notify through REST service for conversation in exp.conversations :
@@
16                   report = device.Trace(model(prompt)) # run inference device.Write(report, exp.conf.outputPath) # results to FS
@@ 4.1. Model Zoo and Evaluation
Model Evaluator. The next step is to evaluate the accuracy degradation of the model due to quantization. To accomplish this, we use MELT’s Model Evaluator component, which is responsible for evaluating the model [^4] We evaluate the non-finetuned variants of the models, as a typical proxy of the accuracy degradation of downstream models. on a given dataset and reporting its accuracy. We leveraged the LLM-Evaluation Harness (Gao et al., 2021) and integrated a custom inference server to serve our converted models from each of the supported backends. This offers a convenient abstraction layer between the frameworks and the evaluation harness. Because of the lack of native support from the frameworks, we had to implement the extraction of token log probabilities to assess the accuracy per downstream dataset [^5] Because of issues with evaluating quantized models on MLC-LLM, we evaluate AWQ (Lin et al., 2023) quantized models with autoawq package as a proxy.. The currently supported datasets are depicted in Table 4 and the results of the evaluation are presented in Sec. 5.4.
@@ 4.2. Automated On-Device Benchmarking
We have automated the build of the framework backends and applications for each platform (e.g., Android, iOS, Linux (CUDA)), along with the conversion binaries for the respective models. We used an M2-powered Mac Studio in the local network to build and package dependencies for mobile targets, especially since Xcode was required to sign app releases on iOS. Specifically, the Android apps were built with Android SDK v.35.0.0 and NDK v.26.1, whereas for iOS we used Xcode 15.2. Installation of packages (.apk and .ipa) was done by the co-ordinator. For the case of JetsonLab, the frameworks and models were compiled on device with CUDA 12.2.
Automator. In order to measure the performance of the respective model on device, we automate the interaction with the chat application. To accomplish this, we use a set of precanned prompts, sampled from the OAAST chat dataset (Köpf et al., 2023), and interact in a multi-turn manner with the LLM. More information about the distribution of these prompts in Sec. 5.2.1.
@@
Framework Backend Version Supported Platforms Quantization MLC-LLM (team, 2023) TVM (Chen et al., 2018) ${}^{\dagger}$ Android (GPU), iOS (Metal), Linux (CUDA) Group Quantization (Shen et al., 2020), GPTQ (Frantar et al., 2022), FasterTransformer Row-wise Quantization llama.cpp (Gerganov, 2023) llama.cpp (Gerganov, 2023) ${}^{\ddagger}$ Android (CPU, GPU), Linux (CUDA) k-quants (llama.cpp Team, 2023) LLMFarm (guinmoon, 2023) llama.cpp (Gerganov, 2023) 7226a8 iOS (Metal) ${}^{\dagger}$ We used version 784530 for supporting Gemma models and Llama-2-7B on Android. ${}^{\ddagger}$ We used version d5ab29 for supporting Gemma models.
@@ 5.1. Experimental Setup
(a) Prefill throughput for MLC-LLM on high-tier devices
(b) Generation throughput for MLC-LLM on high-tier devices
(c) Prefill throughput for llama.cpp on mid-tier devices
(d) Generation throughput for llama.cpp on mid-tier devices
Figure 3. Throughput across frameworks and devices
@@ 5.2. Macro Experiments
(a) MLC-LLM on high-tier devices
(b) llama.cpp on high-tier devices
(c) MLC-LLM on mid-tier devices
(d) llama.cpp on mid-tier devices
@@ 5.2.1. Dataset Qualitative Analysis
(a) MLCChat on Galaxy S23
(b) MLCChat on iPhone 14 Pro
(c) LLMFarm on iPhone 14 Pro
@@ 5.2.2. On-device Runtime
Energy efficiency. Next, we take the same set of models and illustrate the energy discharge (in mAh) per token generated across devices and frameworks in Fig. 4. Overall, we noticed that the trend of larger networks (in terms of parameter size) offering larger discharge rates across devices and frameworks. This is expected as DRAM utilization and memory copies into the SoC registers consume significant energy (Patterson et al., 2022). Notable exceptions to this rule were TinyLlama (3-bit) and Gemma (4-bit), which we aim to investigate with help from upstream maintainers. Last, the CPU execution of llama.cpp offered overall lower energy efficiency, but this could also be attributed to the latency of running inference compared to LLMFarm’s Metal acceleration.

### v4

# angle=0, scale=1, color=black, firstpage=true, position=current page.north, hshift=0pt, vshift=-20pt, contents=

@@ Abstract.
Transformers have recently revolutionized the machine learning (ML) landscape, gradually making their way into everyday tasks and equipping our computers with “sparks of intelligence”. However, their runtime requirements have prevented them from being broadly deployed on mobile. As personal devices become increasingly powerful at the consumer edge and prompt privacy becomes an ever more pressing issue, we explore the current state of mobile execution of Large Language Models (LLMs). To achieve this, we have created our own automation infrastructure, MELT, which supports the headless execution and benchmarking of LLMs on device, supporting different models, devices and frameworks, including Android, iOS and Nvidia Jetson devices. We evaluate popular instruction fine-tuned LLMs and leverage different frameworks to measure their end-to-end and granular performance, tracing their memory and energy requirements along the way.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; November 18–22, 2024; Washington DCprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Lately, however, transformers (Vaswani et al., 2017) have become the go-to architecture for deep learning models, with attention mechanisms offering unparalleled performance and the ability to model long-sequence data with fewer inductive biases. Applied across modalities, including vision, speech and text (Dosovitskiy et al., 2021; Radford et al., 2019, 2023), these models have demonstrated significant performance benefits, across modeling and generation tasks. Another key benefit of these models has been their ability to scale to very large sizes, both in terms of data ingestion and parameter size, without their performance plateauing (Touvron et al., 2023b). This has given birth to “foundation models”, large models that are trained on large corpora of data and act as universal backbones for a series of downstream tasks. For example, Large Language Models (LLMs) (Touvron et al., 2023b; Chung et al., 2022; Hoffmann et al., 2022) have been trained on large parts of the Internet (commoncrawl.org, 2024; Raffel et al., 2020) and are able to tackle downstream tasks without explicit training (Wang et al., 2023; Dong et al., 2023; Vernikos et al., 2023).
@@
To this end, we have created our own infrastructure, named MELT  [^1] Our code can be found at: github.com/brave-experiments/MELT-public., designed to interact, trace and benchmark LLMs across ML frameworks, devices, and ecosystems. With our tool, we automate the interaction with instruction fine-tuned models and capture events and metrics of interest at a granular level, both in terms of performance as well as energy. While in this paper MELT is deployed for LLMs inference, our infrastructure is general enough to support the tracing of any workload, without user intervention. To the best of our knowledge, our tool is the first to support granular on-device energy measurements across device targets (i.e., Android, iOS, Linux) with realistic interactions.
Research Questions. Given the constant developments in mobile and embedded System-On-Chips (SoCs) and the meteoric rise of LLMs, we aspire to measure the deployability of Large Language Models at the consumer edge and identify the bottlenecks that prevent the broad deployment of such workloads. Specifically, the research questions we aim to answer with this study are the following:
@@ 2.1. Transformer Preliminaries
The main contribution of transformers has undoubtedly been the attention mechanism, which captures the relationship between tokens in a sequence from a single source (self-attention) or multiple sources (multi-head attention). Attention is calculated as $A(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V$, where $Q$, $K$, $V$ represent the query, key, and value matrices, respectively, and $d_{k}$ the dimensionality of the key matrix. The inner product boosts closer query-key vectors (relevance), softmax normalizes the dot-product and the multiplication with the value results in the relevant value scores being retrieved. The quadratic complexity of attention, with respect to the sequence length (i.e., the prompt or intermediate tokens), is one of the main bottlenecks of deployment, which has given way to alternatives such as sparse (Beltagy et al., 2020) or approximate (Wang et al., 2020; Ainslie et al., 2023) attention mechanisms, as well as attention-free variants as of lately (Gu and Dao, 2023). Context size refers to the maximal window of tokens a transformer block can pay attention to, whereas the maximum generated length refers to the maximum number of tokens generated as output. Generation ends when an <EOS> (end-of-sequence) token is generated. The auto-regressive nature of decoding means that given an input sequence $X=\{x_{1},x_{2},\dots,x_{t}\}$, the model generates $x_{t+1}$, which is fed to the next generation step. Key-Value cache (Pope et al., 2023) optimizes this by storing intermediary attention states.
@@ 2.2. Large Language Models
What has made Transformers an instant success has been their applicability to various modalities and their scalability to very large parameter sizes without saturating accuracy (Touvron et al., 2023b). This phenomenon has given birth to Foundation Models (FMs), pretrained on huge corpora of data, i.e., text in our case, and act as a great tool for modeling language and a starting point for fine-tuning on downstream tasks. The task of pretraining usually comprises masked or next-word prediction (self-supervised), whereas downstream tasks can include anything from translation to summarization. Instruction fine-tuning (Ouyang et al., 2022) refers to a specific form of fine-tuning where the model is trained on pairs of input-output instructions. Last, alignment is usually the final step of model tuning, typically through reinforcement learning from human (Ouyang et al., 2022) or automated (Bai et al., 2022) feedback, to promote a certain style or content or response that “aligns” with values of the creator (e.g., safety). Training cost generally scales down as we move from pretraining downstream, as do data ingestion needs (Zhou et al., 2024).
@@ 2.3. Current State and Motivating Factors
LLMs democratization. Nevertheless, more and more models offer openly their weights, including models from Meta (Touvron et al., 2023a, b), Mistral AI (Jiang et al., 2023), Google (Google Inc., 2024) and Microsoft (Javaheripi, Mojan and Bubeck, Sébastien, 2024). This creates an excellent opportunity for users to deploy their models locally and even personalize them to their preferences, without data ever leaving their device premises. However, such models remain significantly smaller in scale and still require considerable resources to deploy. Toward this end, new frameworks are emerging for enabling local execution of LLMs across different targets (Gerganov, 2023; team, 2023; Hannun et al., 2023; Mark Sherwood, 2024; Alibaba, 2023; tinygrad, 2023). In this effort, quantization (Shen et al., 2020; Frantar et al., 2022; Lin et al., 2023) is one of the most prominent out-of-the-box solutions for reducing their footprint. Yet another enabler towards this democratization is the broad availability of capable SoCs at cost. Indicatively, from our measurements, a recent M2-based Mac Studio can run Llama-2 (Touvron et al., 2023b) 7B model (4-bit quantized) at a sustained 46.8 tokens/sec.
@@ 3.2. JetsonLab
Pseudocode for MELT experiments. Functionality of undefined methods in comment. Prefixed methods run on the device in prefix (e.g., Monsoon, device).
Input: PhoneLab, JetsonLab, Monsoon, GPIO, YKUSH, device, $Q^{\text{device}}_{\text{experiments}}$, iterations, samplingFrequency, betweenExpSleep
1 PowerOn(device)
2 if device.platform == ”ios” :
3       ConnectBT(device) # connect as HID device via Bluetooth
4       UnlockScreen(device) # unlock screen with passcode over HID
5
6SyncClocks(device) # sync host and guest clocks
7 apiAddress = StartRESTServer() # start REST service on host
8 for exp in $Q^{\text{device}}_{\text{experiments}}$ : # iterate over experiments in the queue
9       Push([exp.model, exp.conversations], device) # push dependencies
10       Apply(exp.conf, exp.model, device) # edit model conf and execution parameters on device
11       for it=0; it¡iterations;++it :
12             StartMonitoring(Monsoon, device)
13             RunExperiment(exp, device)
14             StopMonitoring(Monsoon, device) # disable monitoring
15             CollectMeasurements(exp, device) # get results from FS
16             sleep(betweenExpSleep) # sleep between runs
17
@@ 15             CollectMeasurements(exp, device) # get results from FS
20             GPIO.EnableRail(device.rails) # enable rail through GPIO
21             YKUSH.PowerOn(device) # enable YKUSH USB of device
22             Monsoon.SetVoutCurr(device) # configure Monsoon power out
23             Wait(device) # wait until device is responsive
24
25def StartMonitoring(Monsoon, YKUSH, device):
@@ 23             Wait(device) # wait until device is responsive
27             YKUSH.DisableUSB()
28             Monsoon.MeasurementMode(”on”, samplingFrequency)
29
@@
31             Jetsonlab.ScheduleEvents(samplingFrequency)
32             Jetsonlab.Monitor(”on”) # poll SysFS
33
34def RunExperiment(exp, device, apiAddress):
35       # open app w/ ADB, Bluetooth HID or SSH
36       app = device.OpenApp(exp.backend)
37       Automate(app, model, device) # automate interaction with app
38       http.post(”start”, apiAddress) # notify through REST service
39       for conversation in exp.conversations :
@@ 37       Automate(app, model, device) # automate interaction with app
41                   report = device.Trace(model(prompt)) # run inference
42                   device.Write(report, exp.conf.outputPath) # results to FS
43
@@ 4.1. Model Zoo and Evaluation
Model Evaluator. The next step is to evaluate the accuracy degradation of the model due to quantization. To accomplish this, we use MELT’s Model Evaluator component, which is responsible for evaluating the model [^4] We evaluate the non-finetuned variants of the models, as a typical proxy of the accuracy degradation of downstream models. on a given dataset and reporting its accuracy. We leveraged the LM-Evaluation Harness (Gao et al., 2021) and integrated a custom inference server to serve our converted models from each of the supported backends. This offers a convenient abstraction layer between the frameworks and the evaluation harness. Because of the lack of native support from the frameworks, we had to implement the extraction of token log probabilities to assess the accuracy per downstream dataset [^5] Because of issues with evaluating quantized models on MLC-LLM, we evaluate AWQ (Lin et al., 2023) quantized models with autoawq package as a proxy.. The currently supported datasets are depicted in Table 4 and the results of the evaluation are presented in Sec. 5.4.
@@ 4.2. Automated On-Device Benchmarking
We have automated the build of the framework backends and applications for each platform (e.g., Android, iOS, Linux (CUDA)), along with the conversion binaries for the respective models. We used an M2-powered Mac Studio on the local network to build and package dependencies for mobile targets, especially since Xcode was required to sign app releases on iOS. Specifically, the Android apps were built with Android SDK v.35.0.0 and NDK v.26.1, whereas for iOS we used Xcode 15.2. Installation of packages (.apk and .ipa) was done by the co-ordinator. For the case of JetsonLab, the frameworks and models were compiled on device with CUDA 12.2.
Automator. In order to measure the performance of the respective model on device, we automate the interaction with the chat application. To accomplish this, we use a set of precanned prompts, sampled from the OAAST chat dataset (Köpf et al., 2023), and interact in a multi-turn manner with the LLM. More information on the distribution of these prompts in Sec. 5.2.1.
@@
Framework Backend Version Supported Platforms Quantization MLC-LLM (team, 2023) TVM (Chen et al., 2018) 96a68e Android (GPU), iOS (Metal), Linux (CUDA) Group Quantization (Shen et al., 2020), GPTQ (Frantar et al., 2022), FasterTransformer Row-wise Quantization llama.cpp (Gerganov, 2023) llama.cpp (Gerganov, 2023) b22022 Android (CPU, GPU), Linux (CUDA) k-quants (llama.cpp Team, 2023) LLMFarm (guinmoon, 2023) llama.cpp (Gerganov, 2023) 7226a8 iOS (Metal) We used version 784530 for supporting Gemma models and Llama-2-7B on Android. We used version d5ab29 for supporting Gemma models.    We used version 46bdb4 for supporting Gemma models.
@@ 5.2.2. On-device Runtime
(a) Prefill throughput for MLC-LLM on high-tier devices (GPU)
(b) Generation throughput for MLC-LLM on high-tier devices (GPU)
(c) Prefill throughput for llama.cpp on mid-tier devices (CPU)
(d) Generation throughput for llama.cpp on mid-tier devices (CPU)
Figure 4. Throughput across frameworks and devices
Energy efficiency. Next, we take the same set of models and illustrate the energy discharge (in mAh) per token generated across devices and frameworks in Fig. 5. Overall, we noticed that the trend of larger networks (in terms of parameter size) offering larger discharge rates across devices and frameworks. This is expected as DRAM utilization and memory copies into the SoC registers consume significant energy (Patterson et al., 2022). Notable exceptions to this rule were TinyLlama (3-bit) and Gemma (4-bit), which we aim to investigate with help from upstream maintainers. Last, the CPU execution of llama.cpp offered overall lower energy efficiency, but this could also be attributed to the latency of running inference compared to when using GPU acceleration.
(a) MLC-LLM on high-tier devices (GPU)
(b) llama.cpp on high-tier devices (CPU)
(c) MLC-LLM on mid-tier devices (GPU)
(d) llama.cpp on mid-tier devices (CPU)
(a) MLC-LLM on Galaxy S23 (GPU)
(b) MLC-LLM on iPhone 14 Pro (GPU)
(c) LLMFarm on iPhone 14 Pro (GPU)